# Shortcut Learning in In-Context Learning: A Survey

### v1

@@ 6.1 Differences from Shortcuts in LPLMs
Given that previous reviews have primarily focused on research into LPLMs, we deem it necessary to elucidate the similarities and differences between shortcut learning in LLMs and LPLMs.
@@
The difference between them primarily arises from the distinctions in research paradigms between LPLMs and LLMs. LPLMs typically require fine-tuning for different tasks, whereas LLMs unify various NLU tasks and only need a few samples to guide their inference. Therefore, ICL encounters a broader range of shortcut types compared to those in LPLMs, even though both include shortcut types such as lexicon Niven and Kao (2019), overlap Lai et al. (2021), position Ko et al. (2020), and text style Qi et al. (2021). For instance, since LLMs do not require fine-tuning, the biases inherent in LLMs themselves can strongly influence various tasks, leading to the emergence of the instinctive shortcuts. Furthermore, since LLMs can easily infer answers from few samples, they naturally encounter shortcuts arising from these samples’ group dynamics. The shortcuts related to group dynamics may, to some extent, resemble the challenges faced in long-tailed classification, where models tend to learn answers from head classes rather than tail classes Li et al. (2023a). But besides the skew in categories, LLMs are also influenced by more complex group decision-making process, such as the parity of options Balepur et al. (2024). This can be attributed to LLMs’ stronger comprehension abilities, enabling them to uncover complex correlations that LPLMs are unable to detect.
@@
Another notable difference lies in shortcut mitigation strategies. For LPLMs, a widely adopted mitigation approach involves adjusting the model during fine-tuning, which encompasses techniques such as adversarial training Rashid et al. (2021), regularization Stacey et al. (2022), Product-of-Experts Sanh et al. (2021), reweighting Liu et al. (2021), contrastive learning Song et al. (2023), and other related operations. For LLMs, fine-tuning to mitigate shortcuts is rare, and the primary methods focus on research based on forward propagation. But we can still find extensions of LPLMs in the related research on LLMs. For instance, Calibration methods are similar to Product-of-Experts, both of which adjust the prediction probabilities.
@@ 6.2 Future Directions
More Robust Evaluation Benchmark. Siska et al. (2024) demonstrate that the correlation between the performance exhibited by the model and the prompts is significant, which suggests that the evaluation benchmark itself also exhibits bias. When the model responds to certain similar prompts within the benchmark, it tends to produce similar results. Pacchiardi et al. (2024) also indicate that LLMs can exploit shortcuts in the benchmark to obtain inflated evaluation scores. In addition, data contamination and leakage can also impact the evaluation of LLMs’ shortcuts, as LLMs tend to provide good prediction results directly for benchmarks that have appeared in the training corpus Zhou et al. (2023). Balepur et al. (2024) also imply that leakage of the test set might have an impact on multiple-choice questions without the question. The aforementioned studies suggest that caution should be exercised when interpreting the performance results of LLMs, and the impact of shortcuts should be fully considered. Furthermore, efforts should be made to minimize the potential influence of shortcuts on prediction results before constructing benchmarks.
@@
More Shortcut Related Tasks. Despite the extensive exploration of shortcut learning in ICL, given the diversity of NLP tasks, it remains crucial to expand the boundaries of research tasks. For instance, in Table QA tasks, only a small portion of the entire table is relevant to deriving the answer to a given question, and irrelevant information may become potential shortcuts that affect the prediction outcomes of LLMs Patnaik et al. (2024). These expanded tasks are highly dependent on relevant benchmarks, so designing and constructing corresponding datasets and evaluation methods for new tasks also require attention.
@@
More Interpretability. Although the studies in Section 3 explore the causes of shortcuts from different aspects, most of them are empirical summaries based on experimental observations, lacking an interpretation of the causes of shortcuts from an explainable perspective. For instance, how ICL influences the role of instinctive shortcuts in different NLU tasks, and at which step of the forward propagation in LLMs do these shortcuts exert their impact, is it during aggregation or distribution Wang et al. (2023b)? Can self-explanation methods of LLMs be effective in this process Huang et al. (2023)? Interpretable methods can enhance the intuitive presentation of shortcut learning in ICL and facilitate effective mitigation of shortcuts by addressing their underlying causes.
@@
More Discussion in Shortcut Unknown Scenarios. In the research of LPLMs, shortcut identification is a crucial preliminary step for shortcut mitigation Du et al. (2024), as in many cases, the types of shortcuts in tasks are unknown, or multiple shortcuts coexist. But in ICL, most studies assume that LLMs are vulnerable to shortcuts, or simply presume that the types of shortcuts are known and explore methods to assess performance under these known shortcuts. This idealized and strong assumption of known shortcuts may fail when facing new tasks, leading to limitations in shortcut mitigation solutions. Therefore, exploring how LLMs identify and utilize shortcuts in tasks where the shortcuts are unknown is also one of the subsequent research directions.
@@
Decoupling of Instinctive and Acquired Shortcuts. Based on the discussion in Section 3, instinctive shortcuts are established during the training process of LLMs, while acquired shortcuts originate from demonstrations in the inference process. Instinctive shortcuts should emerge earlier in the lifecycle of LLMs. Consequently, research solely focused on acquired shortcuts may be potentially influenced by instinctive shortcuts, leading to biased conclusions, as discussed by Li et al. (2023b) in LLMs bais. Therefore, there is a need to strengthen the exploration of how these two different types of shortcuts interact with each other. Furthermore, in the research on analyzing and mitigating acquired shortcuts, it is important to consider reducing or eliminating the influence of instinctive shortcuts to ensure the objectivity of the experimental results.

### v2

@@ 6.1 Differences from Shortcuts in LPLMs
Given that previous reviews have primarily focused on research into LPLMs, we deem it necessary to elucidate the similarities and differences between shortcut learning in LLMs and LPLMs to bridge the two different studies.
6.1.1 Shortcut Types
@@
The difference between them primarily arises from the distinctions in research paradigms between LPLMs and LLMs. LPLMs typically require fine-tuning for different tasks, whereas LLMs unify various NLP tasks and only need a few samples to guide their inference. Therefore, ICL encounters a broader range of shortcut types from the few samples compared to those in LPLMs, even though both include shortcut types such as lexicon [102], overlap [20], position [103], and text style [82]. For instance, since ICL do not require fine-tuning, the biases inherent in LLMs themselves can strongly influence various tasks, leading to the emergence of the instinctive shortcuts. Furthermore, since LLMs can easily infer answers from few samples, they naturally encounter shortcuts arising from these samples’ group dynamics. The shortcuts related to group dynamics may, to some extent, resemble the challenges faced in long-tailed classification, where models tend to learn answers from head classes rather than tail classes [104]. But besides the skew in categories, LLMs are also influenced by more complex group decision-making process, such as the parity of options [44]. This can be attributed to LLMs’ stronger comprehension abilities, enabling them to uncover complex correlations that LPLMs are unable to detect with few samples.
6.1.2 Mitigation Strategies
@@
Another notable difference lies in shortcut mitigation strategies. For LPLMs, a widely adopted mitigation approach involves adjusting the model during fine-tuning, which encompasses techniques such as adversarial training [105], regularization [106], Product-of-Experts [107], reweighting [108], contrastive learning [109], and other related operations. These fine-tuning based methods reformulate the parameters of the model to mitigate the model’s focus on shortcuts or samples containing shortcuts. But for LLMs, fine-tuning to mitigate shortcuts is rare, and the primary methods focus on research based on forward propagation. But we can still find extensions of LPLMs in the related research on LLMs. For instance, Calibration methods are similar to Product-of-Experts to some extent, both of which adjust the prediction probabilities. Alternatively, shortcut-based methods are similar to keyword regularization methods in that they both apply specific tokens to a single sample [110]. This intrinsic connection may guide researchers in the direction of transferring mitigation methods from LPLMs to ICLs.
6.1.3 Shortcut Search
Based on the apriorism premise that the shortcut is known, it is easy to make actions to alleviate the shortcut for LPLMs. However, this requires expert knowledge, and its generalization to complex and variable applications is limited [111]. Therefore, some methods explore scenarios where shortcuts are unknown and search for biased samples and shortcuts using knowledge distillation [112] and interpretable methods, such as attention scores [113], shortcut degree [114], saliency technique [115], integrated gradients [96, 116], counterfactual attribution [117], etc. These methods may also be used in the study of shortcut learning in ICL [50]. For example, Satyapriya et al. determine the facilitating effect of post hoc explanations on LLMs [118]. Besides, interpretable methods also help to search for higher quality samples, which again helps to alleviate shortcuts [119, 120]. Still, interpretable methods in shortcut learning still need to be further explored, as described in Section 6.2.
@@ 6.2.1 More Robust Evaluation Benchmark
Studies have demonstrated that the correlation between the performance exhibited by the model and the prompts is significant, which suggests that the evaluation benchmark itself also exhibits bias[121]. When the model responds to certain similar prompts within the benchmark, it tends to produce similar results. Lorenzo et al. also indicate that LLMs can exploit shortcuts in the benchmark to obtain inflated evaluation scores [60]. In addition, data contamination and leakage can also impact the evaluation of LLMs’ shortcuts, as LLMs tend to provide good prediction results directly for benchmarks that have appeared in the training corpus [122]. Nishant et al. also imply that leakage of the test set might have an impact on multiple-choice questions without the question [44]. The aforementioned studies suggest that caution should be exercised when interpreting the performance results of LLMs, and the impact of shortcuts should be fully considered. Furthermore, efforts should be made to minimize the potential influence of shortcuts on prediction results before constructing benchmarks.
6.2.2 More Shortcut Related Tasks
@@
Despite the extensive exploration of shortcut learning in ICL, given the diversity of NLP tasks, it remains crucial to expand the boundaries of research coverage and tasks. As shown in Table 1, the performance of some shortcut types on some tasks has not been fully explored and may contain new discoveries. Besides, on more tasks such as Table QA, only a small portion of the entire table is relevant to deriving the answer to a given question, and irrelevant information may become potential shortcuts that affect the prediction outcomes of LLMs [123]. For another example, in the planing task, which determines the process of the sequence of actions needed to achieve the goal, LLMs have also been observed to focus on short-horizon and low-level planning, which may also become a shortcut learning [124]. These expanded tasks are highly dependent on relevant benchmarks, so designing and constructing corresponding datasets and evaluation methods for new tasks also require attention.
6.2.3 More Interpretability
@@
Although the studies in Section 3 explore the causes of shortcuts from different aspects, most of them are empirical summaries based on experimental observations, lacking an interpretation of the causes of shortcuts from an explainable perspective. For instance, how ICL influences the role of instinctive shortcuts in different NLP tasks, and at which step of the forward propagation in LLMs do these shortcuts exert their impact, is it during aggregation or distribution [125]? Interpretable methods can enhance the intuitive presentation of shortcut learning in ICL and facilitate effective mitigation of shortcuts by addressing their underlying causes.
However, there are still many limitations in applying these interpretable methods on LLMs. For example, the commonly used integrate gradient methods requires LLMs to perturb and assign values to each feature in the input. The time cost is unacceptable on large-scale datasets and LLMs. Similarly, the saliency technique requires performing at least one backpropagation on the LLMs to compute the saliency score for each element, which is not applicable for black-box LLMs [125]. Therefore, the time cost and the opacity of black-box LLMs still need to be considered in the interpretable study of shortcut learning. A feasible approach is to attribute LLMs through small-scale models [118], but this ignores the differences between small-scale models and LLMs [126]. Distilling knowledge from LLMs into the small-scale models could be a potential solution [127]. Another viable way is to reduce the complexity of interpretable evaluation through heuristic search methods, but this may also lead to suboptimal results [128]. For black-box LLMs, the self-explanatory ability of LLMs can be a shortcut learning exploration tool in this process [87]. All in all, interpretable analysis in shortcut learning still needs to be further explored in ICL.
6.2.4 More Discussion in Shortcut Unknown Scenarios
@@
In the research of LPLMs, shortcut identification is a crucial preliminary step for shortcut mitigation [10], as in many cases, the types of shortcuts in tasks are unknown, or multiple shortcuts coexist. But in ICL, most studies assume that LLMs are vulnerable to shortcuts, or simply presume that the types of shortcuts are known and explore methods to assess performance under these known shortcuts. This idealized and strong assumption of known shortcuts may fail when facing new tasks, leading to limitations in shortcut mitigation solutions. Therefore, exploring how LLMs identify and utilize shortcuts in tasks where the shortcuts are unknown is also one of the subsequent research directions. In this scenario, interpretable methods may be used as a pre-technical means for shortcut mining.
6.2.5 Decoupling of Instinctive and Acquired Shortcuts
@@
Based on the discussion in Section 3, instinctive shortcuts are established during the training process of LLMs, while acquired shortcuts originate from demonstrations in the inference process. Instinctive shortcuts should emerge earlier in the lifecycle of LLMs. Consequently, research solely focused on acquired shortcuts may be potentially influenced by instinctive shortcuts, leading to biased conclusions. Jannik et al. show that label relationships inferred from pretraining appear to have permanent effects that can not be overcome by observations in context [37]. When the LLMs are unfamiliar with the task, it is difficult to solve the problem by recalling the pre-trained knowledge. In this case, only similar examples can improve model performance by providing more direct shortcuts [129]. Moreover, in other application scenarios, such as Retrieval Augmented Generation, LLMs has a strong bias towards leveraging contextual information to answer questions, while minimally relying on internal knowledge [130], again suggesting that LLMS may be more affected by acquired shortcuts. Therefore, there is a need to strengthen the exploration of how these two different types of shortcuts interact with each other. Furthermore, in the research on analyzing and mitigating acquired shortcuts, it is important to consider reducing or eliminating the influence of instinctive shortcuts to ensure the objectivity of the experimental results.
6.2.6 Exploration of Multiple Shortcut Coexistence
This can be regarded as an extension of the decoupling of instinctive and acquired shortcuts. Li et al. show that mitigating one shortcut increases the dependence on other shortcuts [131]. So, when multiple shortcuts coexist, quantifying the mutual influence between shortcuts and giving the corresponding mitigation plan is also one of the future research directions. Some technical gains of Intersectional Debias can guide the corresponding research methods, such as nullspace projection [132]. But in complex scenarios, when the given shortcut is unknown, the corresponding research faces more challenges.
@@ 7 Conclusion
This work was supported by China Postdoctoral Science Foundation Funded Project (No. 2024M761122), the National Natural Science Foundation of China (No.62476111), and the Department of Science and Technology of Jilin Province, China (No.20230201086GX)